There are 435 congressmen I the united state. Of those 435, 24 of them are from Arizona, oregon, and Washington. Arizona has five more congressmen than oregon Washington has twice as many as oregon does. how many congressmen does oregon have?
Answer by Boreal(15235)

The question is wrong.
or=x
az=x+5
wa=2(x+5)
They add to 24
x+x+5+2x=24
4x+5=24
4x=19, and this doesn't give an integer.
Arizona has 9, Washington 10, and Oregon 5 (I live in the NW and I lived 37 years in Arizona.)